I need help with this problem please. I'm getting the wrong numbers. Thank You!

Exercise 4-4 Your answer is partially correct. Try again. On October 5, 2017, Diamond in the Grouper Group Inc.s board of directors decided to dispose of the Blue Division. A formal plan was approved Diamond derives approximately 79% of its income from its human resources management practice. The Blue Division gets contracts to perform human resources management on an outsourced basis. The board decided to dispose of the division because of unfavourable operating results Net income for Diamond was $91,400 for the fiscal year ended December 31, 2017 (after a charge for tax at 30% and after a writedown for the Blue assets). Income from operations of the Blue Division accounted for $4,600 (after tax) of this amount. Because of the unfavourable results and the extreme competition, the board believes that it cannot sell the business intact. Its final decision is to auction off the office equipment. The equipment is the divisions only asset and has a carrying value of $27,000 at October 5, 2017. The board believes that proceeds from the sale will be approximately $5,000 after the auction expenses. Currently, the estimated fair value of the equipment is $8,000. The Blue Division qualifies for treatment as a discontinued operation. Diamond prepares financial statements in accordance with ASPE. Prepare a partial income statement for Diamond in the Grouper Group. The income statement should begin with income from continuing operations before income tax. (Enter negaive amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)Prepare a partial income statement for Diamond in the Grouper Group. The income statement should begin with income from continuing operations before income tax. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).) Diamond in the Grouper Group Inc. Partial Income Statement For the Year Ended December 31, 2017 Income from Continuing Operations before Income Tax 139958.6 Income Tax 41988 Income from Continuing Operations Discontinued operations Income from Operations of the Discontinued Blue Division, less Applicable Income Tax 97971 4600 Loss from impairment of assets of Discontinued operations, less Applicable Income Tax Recovery -15429 -10829 Net Income / (Loss) 87142
Answer #1

Answer:

Diamond in the Grouper Group Inc.

Partial Income Statement

For the year ended December 31, 2017

Income from continuing operations before income tax

$146,000
Less: Income tax (43,800)
Income from continuing operations 102,200
Discontinued operations
Income from operations of the discontinued Blue Division, less applicable income tax 4,600
Loss from impairment of assets of discontinued operations, less applicable income tax recovery (15,400)
(10,800)
Net Income / (Loss) $91,400

Working Notes :

1. Income from continuing operations -

  • Income from Diamond - $(91,400 ÷ 70) × 100 = $130,571.43
  • Blue division asset written off = $27,000 - 5,000 = $22,000
  • Blue division's income = $( 4,600÷70)×100 = $6571.43

Income from continuing operations = $ (130,571.43 + 22,000 - 6571.43) = $ 146,000

2. Loss from impairment of assets, less applicable income tax recovery = (27,000 - 5,000) - 30% = $15,400